How do you do problems like 2x-3y=-14 ans 3x+4y=-4 using elimination?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Multiply the first equation by 4 and the second equation by 3. The resulting coefficients on will be additive inverses. Then, when you add the two equations term-by-term, the term will be eliminated (hence the name of the method) and you will be left with one equation in the single unknown which can then be solved by ordinary methods. Once you have determined a value for , you can substitute that value for in either of the original equations and then solve for

Additional information. Stan's solution below inadvertently has the values of the x and y coordinates reversed. I received the e-mail complaining about it. Here is my response:
I didn't give you an answer, so how could it be backwards? If you are responding to Stan's answer, take it up with him. I know Stan, he is a really good guy. If he does a problem incorrectly, he ALWAYS refunds 10 times the amount he received for his services. I do the same thing. We can do that because 10 times nothing is nothing. Now quit whining and go do your homework.


John


Answer by stanbon(75887)

How do you do problems like
2x-3y=-14
3x+4y=-4 using elimination?
---------
Multiply the 1st Eq by 3; Multiply the 2nd Eq. by 2
6x-9y = -42
6x+8y = -8
-------------
Subtract the 1st from the 2nd:
17y = 34
y = 2
-----
Substitute to solve for "x":
2x - 3*2 = -14
2x = -8
x = -4
---------------------------
Solution: (2,-4)
=================================
Cheers,
Stan H.